to get out of jail free in the board game monoloply you have to roll doubles with a pair of standard dice. Determine odds in favour of getting out of jail on first or second roll. the answer is 11:25, show work
Answer #1

as threre are 6 posssible outcome of doubles out of 26 total outcome ;

hence probability of a double =6/36 =1/6

P(getting doubles on first or second roll)=P(double on first roll)+P(not on first roll but on second roll)

=(1/6)+(5/6)*(1/6)=11/36

therefore odds in favour of getting out of jail on first or second roll =p/(1-p)=(11/36)/(1-11/36)=(11/36)/(25/36)=11/25 =11:25